OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5 -3x-9 = -3x-3 -5 continue